### v1

# Machine-Learning-Enhanced Optimization of Noise-Resilient Variational Quantum Eigensolvers

Abstract
Variational Quantum Eigensolvers (VQEs) are a powerful class of hybrid quantum-classical algorithms designed to approximate the ground state of a quantum system described by its Hamiltonian. VQEs hold promise for various applications, including lattice field theory. However, the inherent noise of Noisy Intermediate-Scale Quantum (NISQ) devices poses a significant challenge for running VQEs as these algorithms are particularly susceptible to noise, e.g., measurement shot noise and hardware noise.
In a recent work, it was proposed to enhance the classical optimization of VQEs with Gaussian Processes (GPs) and Bayesian Optimization, as these machine-learning techniques are well-suited for handling noisy data. In these proceedings, we provide additional insights into this new algorithm and present further numerical experiments. In particular, we examine the impact of hardware noise and error mitigation on the algorithm’s performance. We validate the algorithm using classical simulations of quantum hardware, including hardware noise benchmarks, which have not been considered in previous works. Our numerical experiments demonstrate that GP-enhanced algorithms can outperform state-of-the-art baselines, laying the foundation for future research on deploying these techniques to real quantum hardware and lattice field theory setups.
1 Introduction
Quantum hardware with several hundreds of qubits can already be harnessed to outperform classical computers on specific tasks (see, e.g., Ref. [1]). These tasks, however, currently have no practical applications and are specifically designed to be challenging for classical computers. As we have entered the Noisy Intermediate-Scale Quantum (NISQ) era of quantum computing, the primary challenge is to develop algorithms for NISQ devices that may demonstrate quantum advantage in tasks of practical relevance, such as quantum chemistry [2] or quantum field theories [3, 4].
Particularly relevant for NISQ devices are hybrid quantum-classical algorithms, such as the Variational Quantum Eigensolver (VQE) [5, 6]. VQEs can approximate the ground state of quantum Hamiltonians by classical variational optimization of a parametrized quantum circuit. The quantum computer measures the energy of the quantum system for a given set of parameters, i.e., the objective of the minimization problem, and a classical optimization routine is used to find a better set of parameters. Efficient exploration of the parameter space is crucial for successful optimization.
Nakanishi et al. [7] demonstrated that, under certain conditions, the functional form of the VQE’s objective function can be explicitly derived and utilized to develop a more efficient optimization routine. The resulting Nakanishi-Fuji-Todo (NFT) algorithm [7] uses two measurements to sequentially optimize one parameter of the quantum circuit at a time, by fitting the objective in a one-dimensional subspace. Indeed, this constitutes a sequential minimal optimization (SMO) [8], which differs from global optimization protocols such as gradient-based approaches.
Despite its efficiency, the NFT algorithm faces several challenges, primarily due to the noise of current quantum hardware. In particular, the algorithm relies on noisy measurements to identify the global minimum in the one-dimensional subspace, which often hinders convergence and complicates optimization. Machine learning offers a promising approach to addressing the challenge of noisy measurements and enhancing the NFT algorithm with more flexibility. In a recent work [9], Nicoli et al. proposed an optimization algorithm called NFT-with-EMICoRe (EMICoRe for short), which extends the NFT framework by incorporating Gaussian Process Regression (GPR) and Bayesian Optimization (BO) for a data-driven and potentially more noise-resilient sequential minimal optimization scheme. However, this work [9] focused solely on measurement shot noise, neglecting the impact of quantum hardware noise, such as decoherence, depolarization, and crosstalk errors.
In these proceedings, we build on prior works [7, 9] and demonstrate that EMICoRe outperforms the NFT baseline, even in the presence of hardware noise. We start in Sec. 2 by reviewing the theoretical foundations of VQE, Gaussian Processes, Bayesian Optimization. In Sec. 3, we introduce the EMICoRe algorithm. This is followed by numerical experiments in Sec. 4, which involve classical simulations of noisy quantum hardware, comparing the results with and without error mitigation. We conclude and provide an outlook in Sec. 5.
2 Theoretical Foundations
2.1 Variational Quantum Eigensolver (VQE)
Computing the ground state energy of a quantum system described by a Hamiltonian $H$ and the corresponding ground state wave function |$\ket{\psi_{\mathrm{GS}}}$ is a key challenge in many fields of physics. The Hamiltonian of an $N$-qubit system can be represented as
$H=\sum_{\alpha=1}^{T}h_{\alpha}P_{\alpha},$ (1)
where $P_{\alpha}\in\{X,Y,Z,I\}^{\otimes N}$ is a tensor product of single-qubit Pauli operators $X,Y,Z$ and the identity operator $I$. The $h_{\alpha}$ are real coefficients, and $T$ is the number of individual Pauli terms.
In general, it is challenging to compute the ground state wave function |$\ket{\psi_{\mathrm{GS}}}$ because it requires diagonalizing the Hamiltonian $H$, which is a $\mathbb{C}^{2^{N}\times 2^{N}}$-dimensional matrix for an $N$-qubit system and thus scales exponentially with the number of qubits. An alternative approach to exact diagonalization is to compute a variational approximation of the ground state. This can be achieved, for instance, by using the Variational Quantum Eigensolver (VQE) [5, 6], where a parametric quantum state $|\psi(\boldsymbol{\theta})\rangle$ is prepared on a quantum computer. This quantum state, which depends on $D$ angular parameters $\boldsymbol{\theta}\in[0,2\pi)^{D}$, is obtained through a sequence of $D^{\prime}$ ($\geq D$) unitary quantum gate operations $U(\boldsymbol{\theta})=U_{1}\cdots U_{D^{\prime}}$ acting on an initial state $|\psi_{0}\rangle$. A subset $D$ of these operations are Pauli rotation gates, $U_{d^{\prime}(d)}(\theta_{d})=R_{P_{d}}(\theta_{d})=\exp(-i\theta_{d}P_{d})$, where $P_{d}\in\{X,Y,Z\}$ are the Pauli matrices and $\theta_{d}$ are the angular parameters. The full variational state can thus be written as $|\psi(\boldsymbol{\theta})\rangle=U(\boldsymbol{\theta})|\psi_{0}\rangle$. After the preparation of the quantum state $|\psi(\boldsymbol{\theta})\rangle$ on the quantum computer, the state is measured in the basis of the individual Pauli strings to obtain the corresponding measured energy given by
$\tilde{E}(\boldsymbol{\theta})$ $=E^{*}(\boldsymbol{\theta})+\varepsilon,\quad$ (2)
$\textrm{where}\quad E^{*}(\boldsymbol{\theta})$ $=\left\langle\psi(\boldsymbol{\theta})|H|\psi(\boldsymbol{\theta})\right\rangle=\left\langle\psi_{0}\left|U(\boldsymbol{\theta})^{\dagger}HU(\boldsymbol{\theta})\right|\psi_{0}\right\rangle\text{. }$ (3)
The energy $E^{\ast}(\boldsymbol{\theta})$ is the true energy that that would be obtained if the measurement on the quantum computer were error-free. Typically, a quantum computer experiences both Gaussian noise, arising from the statistical nature of repeated measurements (i.e., measurement shot noise), and hardware noise, which results from imperfections in qubits, gates, and measurements [10]. Both types of noise are encoded in the noise term $\varepsilon$ in Eq. 2.
As shown in the green box of Fig. 1, the VQE protocol can be divided into the following steps:
  1. 1.
    Initialize a random quantum state with random parameters $\boldsymbol{\theta}$.
  2. 2.
    Measure the expectation value $\tilde{E}(\boldsymbol{\theta})=\left\langle\psi(\boldsymbol{\theta})|H|\psi(\boldsymbol{\theta})\right\rangle+\varepsilon$ on the quantum computer.
  3. 3.
    Use a classical non-linear optimizer, such as gradient-based or gradient-free methods, to determine new parameters ${\boldsymbol{\hat{\theta}}}$ that minimize $\tilde{E}(\boldsymbol{\theta})$.
  4. 4.
    Iterate to step 2 until $\tilde{E}(\boldsymbol{\theta})$ converges to the ground state energy.
Figure 1: Illustration of the VQE workflow. For more details, see the text.
Under mild assumptions for the above described variational circuit, Nakanishi et al. derived an analytical expression of the VQE’s objective functional [7],
$E^{\ast}(\boldsymbol{\theta})=\boldsymbol{b}^{\top}\cdot\left[\bigotimes^{D}_{d=1}\left(\begin{array}[]{c}\cos{(\theta_{d})}\\ \sin{(\theta_{d})}\\ 1\end{array}\right)\right]\textrm{ , }\quad\forall\boldsymbol{\theta}\in[0,2\pi)^{D},$ (4)
where $\boldsymbol{b}$ is a set of arbitrary coefficients and $\theta_{d}$ is a $D-$dimensional vector of angular parameters for the variational quantum circuit. In particular, this functional form allows the following observation: when keeping all but one parameter constant, i.e.,
$\boldsymbol{\theta}^{(t)}_{d}=(\theta_{1},\ldots,\theta_{d-1},\theta_{d},\theta_{d+1},\ldots,\theta_{D})^{\top}\quad\mathrm{with}\quad\theta_{i\neq d}=\textrm{const.,}$., (5)
the form of the objective function for this one-dimensional subspace of the whole optimization problem reduces to
$E^{\ast(t)}(\boldsymbol{\theta}^{(t)}_{d})$ $=\sum_{k=1}^{K}h_{k}\left\langle\psi_{k}\left|U_{d}^{\dagger}\left(\theta^{(t)}_{d}\right)P_{k}U_{d}\left(\theta^{(t)}_{d}\right)\right|\psi_{k}\right\rangle$ (6)
$=a^{(t)}_{1,d}\cos\left(\theta^{(t)}_{d}-a^{(t)}_{2,d}\right)+a^{(t)}_{3,d},$
where $a^{(t)}_{\ell,d}(\ell=1,2,3)$ denote parameters that are independent of $\theta^{(t)}_{d}$, while the superscript $(t)$ refers to the optimization step. As a result, it is enough to measure three points in the one-dimensional space to fit a function over the entire subspace and analytically determine the minimum.
The NFT optimization algorithm [7] uses this property and sequentially optimizes each parameter in the set $\boldsymbol{\theta}$ by requiring (at least) two measurements for each optimization step. The new observed points are deterministically chosen along the axis $d$, i.e.,
$\boldsymbol{\Theta}^{\prime}=\left(\boldsymbol{\theta}_{1}^{\prime},\boldsymbol{\theta}_{2}^{\prime}\right)=\left\{{\boldsymbol{\hat{\theta}}}^{(t-1)}-\alpha\boldsymbol{e}_{d},{\boldsymbol{\hat{\theta}}}^{(t-1)}+\alpha\boldsymbol{e}_{d}\right\}\,,$ (7)
where $\boldsymbol{\hat{\theta}}^{(t-1)}$ denotes the previous best point from step $(t-1)$ [^1] The original paper from Nakanishi et al. [7] proposed a shift $\alpha$ of $\frac{\pi}{2}$ for Eq. 7. In the work by Nicoli et al. [9] it was suggested, that equidistant points for probing the subspace should be used, i.e., a shift of $\frac{2\pi}{3}$ should be preferred. Later, it was derived by Anders et al. [11] that the offset of $\frac{2\pi}{3}$ provides a constant variance over the whole 1D-subspace. In a private communication with the authors, it was reported that, because of a bug, the actual value used in the experiments of [9] was $\frac{\pi}{3}$ instead of $\frac{2\pi}{3}$ as reported in the paper. Although this did not lead to significant differences in the numerical experiments, in order to be consistent with the prior work, we will stick to the convention of using $\frac{\pi}{3}$ for NFT.. This approach is advantageous because it enables analytical optimization of each parameter within its respective one-dimensional subspace, eliminating the need to compute gradients as required by gradient-based methods, such as SPSA [12, 13]. Furthermore, gradient-based approaches may optimize all parameters simultaneously during each gradient update but do not guarantee analytical minimization.
2.2 Gaussian Processes and Bayesian Optimization
The NFT algorithm can find the exact analytical minimum within the one-dimensional subspace if and only the measurements on the quantum computer are exact, i.e., an infinite number of shots and no hardware noise. However, this is not a real-world scenario. In fact, due to the unavoidable hardware noise on current NISQ hardware, and the inherent statistical uncertainty from repeated measurements (shots), it may be beneficial to enhance existing algorithms with frameworks like Bayesian Optimization (BO) [14] and Gaussian Process Regression (GPR) [15]. These techniques, widely established in machine learning, are often well-suited to deal with noisy data [15]. Let $E^{\ast}(\cdot):\mathcal{X}\mapsto\mathbb{R}$ be an unknown (black-box) objective function to be minimized (i.e., the energy of a VQE in our case). With the aid of a function that suggests promising new points to observe, often referred to as acquisition function [14], one iteratively approximates the target objective with a surrogate function $E(\boldsymbol{\theta})\approx E^{\ast}(\boldsymbol{\theta})$, sampled from the GP. At each optimization step, after new points are measured, those are added to the previous set of measured points, used to update the GP.
A common choice as a surrogate model is a GP regression model with a one-dimensional Gaussian likelihood and GP prior [9]
$p(\tilde{E}\mid\boldsymbol{\theta},E(\cdot))=\mathcal{N}_{1}\left(\tilde{E};E(\boldsymbol{\theta}),\sigma^{2}\right),\quad p(E(\cdot))=\operatorname{GP}(E(\cdot);\nu(\cdot),k(\cdot,\cdot)).$ (8)
Here, $\sigma^{2}$ is the variance of the measurement noise $\varepsilon$, and the prior mean function $\nu(\cdot)$ is set to 00 in this work. In the case of VQEs a zero-mean prior is a reasonable choice because the target function $E^{\ast}(\boldsymbol{\theta})$ one aims to model is a linear combination of functions centered around zero (up to a potential constant offset which stems from Eq. 4). The function $k(\cdot,\cdot)$ is the kernel (covariance) function encoding the expected correlation between two measured points. We refer to  Section 3.1 for more details on the kernel function we use in this proceeding. The GP is trained on noisy measurements $\tilde{E}=E^{\ast}(\boldsymbol{\theta})+\varepsilon$ that were collected throughout previous optimization steps.
Let $\{\boldsymbol{\Theta},\boldsymbol{\tilde{E}}\}$ be $N$ training samples, where $\boldsymbol{\Theta}=(\boldsymbol{\theta}_{1},\ldots,\boldsymbol{\theta}_{N})\in\mathcal{X}^{N}$ are the angular inputs of the parametrized quantum circuit, and $\boldsymbol{\tilde{E}}=(\tilde{E}_{1},\ldots,\tilde{E}_{N})^{\top}\in\mathbb{R}^{N}$ are the energies measured for the corresponding angular parameters $\boldsymbol{\theta}_{i}\in\boldsymbol{\Theta}$. The posterior of a GP regression model, see Eq. 8, is still a GP, i.e., $p(E(\cdot)\mid\boldsymbol{\Theta},\boldsymbol{\tilde{E}})=\operatorname{GP}\left(E(\cdot);\mu_{\boldsymbol{\Theta}}(\cdot),s_{\boldsymbol{\Theta}}(\cdot,\cdot)\right)$. Thus, in Ref. [9] it was shown that for arbitrary $M$ test points $\boldsymbol{\Theta}^{\prime}=\left(\boldsymbol{\theta}_{1}^{\prime},\ldots,\boldsymbol{\theta}^{\prime}{}_{M}\right)\in\mathcal{X}^{M}$, the posterior, or predictive distribution of the function values $\boldsymbol{E}^{\prime}=\left(E\left(\boldsymbol{\theta}_{1}^{\prime}\right),\ldots,E\left(\boldsymbol{\theta}_{M}^{\prime}\right)\right)^{\top}\in\mathbb{R}^{M}$ is the $M$-dimensional Gaussian with its mean and covariance analytically given
$\begin{gathered}p\left(\boldsymbol{E}^{\prime}\mid\boldsymbol{\Theta},\boldsymbol{\tilde{E}}\right)=\mathcal{N}_{M}\left(\boldsymbol{E}^{\prime};\boldsymbol{\mu}_{\boldsymbol{\Theta}}^{\prime},\boldsymbol{S}_{\boldsymbol{\Theta}}^{\prime}\right),\quad\text{ where }\\ \boldsymbol{\mu}_{\boldsymbol{\Theta}}^{\prime}=\boldsymbol{K}^{\prime\top}\left(\boldsymbol{K}+\sigma^{2}\boldsymbol{I}_{N}\right)^{-1}\boldsymbol{\tilde{E}},\quad\boldsymbol{S}_{\boldsymbol{\Theta}}^{\prime}=\boldsymbol{K}^{\prime\prime}-\boldsymbol{K}^{\prime\top}\left(\boldsymbol{K}+\sigma^{2}\boldsymbol{I}_{N}\right)^{-1}\boldsymbol{K}^{\prime}.\end{gathered}$ (9)
The notation with a prime refers to new, unseen, points with the corresponding energies predicted by the GP. Here, $\boldsymbol{K}=k(\boldsymbol{\Theta},\boldsymbol{\Theta})\in\mathbb{R}^{N\times N},\boldsymbol{K}^{\prime}=k\left(\boldsymbol{\Theta},\boldsymbol{\Theta}^{\prime}\right)\in\mathbb{R}^{N\times M}$, and $\boldsymbol{K}^{\prime\prime}=k\left(\boldsymbol{\Theta}^{\prime},\boldsymbol{\Theta}^{\prime}\right)\in\mathbb{R}^{M\times M}$ are the train, train-test, and test kernel matrices, respectively, where $k\left(\boldsymbol{\Theta},\boldsymbol{\Theta}^{\prime}\right)$ denotes the kernel matrix evaluated at each column of $\boldsymbol{\Theta}$ and $\boldsymbol{\Theta}^{\prime}$ such that $\left(k\left(\boldsymbol{\Theta},\boldsymbol{\Theta}^{\prime}\right)\right)_{n,m}=k\left(\boldsymbol{\theta}_{n},\boldsymbol{\theta}_{m}\right)$. Moreover, $\boldsymbol{I}_{N}\in\mathbb{R}^{N\times N}$ denotes the identity matrix. The subscript $\boldsymbol{\Theta}$ in Eq. 9 denotes which input points were used to train the GP. Note that the predictive mean and covariance are given in their vector or matrix form, respectively. The predictive mean can be computed as a function of one arbitrary input vector $\mu_{\boldsymbol{\Theta}}(\boldsymbol{\theta}^{\prime})=\boldsymbol{K}^{\prime\top}\left(\boldsymbol{K}+\sigma^{2}\boldsymbol{I}_{N}\right)^{-1}\boldsymbol{\tilde{E}}$ with $\boldsymbol{K}^{\prime}=k\left(\boldsymbol{\Theta},\boldsymbol{\theta}^{\prime}\right)\in\mathbb{R}^{N}$. The computation of a single predictive covariance $s_{\boldsymbol{\Theta}}(\cdot,\cdot)$ follows analogously. For more details on how to compute the predictive mean and variance we refer to Ref. [9].
The knowledge gathered by the GP during optimization, as the number of measurements increases, can not only be exploited for regression but can also be leveraged to choose the next points to observe. This is achieved by combining GPR and BO and solving the following maximization problem
$\max_{\boldsymbol{\Theta^{\prime}}}{a_{\boldsymbol{\Theta}^{(t-1)}}(\boldsymbol{\Theta^{\prime}})}\,$
where $a_{\boldsymbol{\Theta}^{(t-1)}}(\cdot)$ is the so-called acquisition function. This latter uses a GP model trained on previously observed data pairs $\{\boldsymbol{\Theta}^{(t-1)},\boldsymbol{\tilde{E}}\}$, collected over the prior $(t-1)$ iterations, to assess how promising a set of new input points $\boldsymbol{\Theta^{\prime}}$ may be. Points that maximize the acquisition function are likely to be the most informative points to measure at the next iteration, such that the energy is minimized.
3 Method
In this section, we leverage the functional form of the VQE objective as derived in [7] and we construct a physics-informed kernel suited for GPR. Furthermore, we recall an innovative type of acquisition function and the concept of confident region (CoRe) as introduced in [9]. Finally, in Section 3.2.1, we show how the NFT baseline algorithm can be improved to achieve a more flexible and noise-resilient algorithm when combined with BO and GPR.
3.1 VQE Kernel
A GP needs to be equipped with a kernel (covariance) function. While there are many types of such functions, one popular example is the Gaussian-RBF kernel [15], which encodes into the GP the belief that input vectors that are close in the input space strongly correlate compared to points that are far apart, with the correlation strength decaying exponentially with the distance. Although this might be a reasonable assumption, an exponential drop in correlation, like in the Gaussian-RBF kernel, may not be representative of the underlying features of every problem. For this reason, wisely choosing the kernel for a GP regression task can substantially affect the performance. In Ref. [9] the authors proposed the so-called VQE kernel:
$k^{\mathrm{VQE}}\left(\boldsymbol{\theta},\boldsymbol{\theta}^{\prime}\right)=\sigma_{0}^{2}\prod_{d=1}^{D}\left(\frac{\gamma^{2}+\cos\left(\theta_{d}-\theta_{d}^{\prime}\right)}{1+\gamma^{2}}\right),$ (10)
where $D$ is the number of angular parameters of the parametrized quantum circuit, $\sigma_{0}$ corresponds to the prior variance, and $\gamma^{2}\geq 1$ controls the smoothness of the kernel. Both $\sigma_{0}$ and $\gamma^{2}$ are hyperparameters of the kernel.
It was shown that the VQE kernel can be decomposed as
$k^{\mathrm{VQE}}\left(\boldsymbol{\theta},\boldsymbol{\theta}^{\prime}\right)=\boldsymbol{\phi}(\boldsymbol{\theta})^{\top}\boldsymbol{\phi}(\boldsymbol{\theta}^{\prime}),\quad\textrm{with}\quad\boldsymbol{\phi}(\boldsymbol{\theta})=\sigma_{0}\left(1+\gamma^{2}\right)^{-D/2}\left[\otimes_{d=1}^{D}\left(\gamma,\cos\theta_{d},\sin\theta_{d}\right)^{\top}\right]$ (11)
where the feature maps $\boldsymbol{\phi}(\boldsymbol{\theta})$ fulfill the same VQE functional form derived in Eq. 4. Since a GP regression model is only defined in terms of inner products of the feature maps, one can replace all of these occurrences with the kernel described above – a procedure often known as kernel trick [15].
Using the VQE kernel function instead of, e.g., the RBF-kernel, ensures that the GP samples functions that match the required functional form from Eq. 4. Consequently, as described in Eq. 6, when optimizing on a one-dimensional subspace, i.e., changing only one parameter at a time, three (noiseless) observed points allow to uniquely identify the target function [7] over the entire one-dimensional subspace. However, in the presence of statistical uncertainty due to repeated measurements (shots), the target function in the subspace can be identified up to some statistical uncertainty. Therefore, the GP will be able to provide energy estimates, and their corresponding uncertainties, for new measurements in the subspace.
Furthermore, it has been shown that the decomposition of the VQE kernel in Eq. 11 can be generalized so that each entry of $\boldsymbol{\theta}$ may parametrize multiple gates at once [9]. This property becomes crucial if the Hamiltonian is invariant under certain symmetries, which results in multiple parameters being identical and requiring simultaneous rotation.
3.2 EMICoRe: Expected Maximum Improvement over Confident Regions
To perform Bayesian Optimization (BO) in combination with GPR, one should pick a suitable acquisition function to identify the most promising points to measure at each optimization step. A novel acquisition function termed Expected Maximum Improvement over Confident Regions (EMICoRe) was introduced in the context of GPR for VQEs [9]. This acquisition function relies on the idea of Confident Region (CoRe) mathematically defined as
$\mathcal{Z}_{\boldsymbol{\Theta}}=\left\{\boldsymbol{\theta}\in\mathcal{X};s_{\boldsymbol{\Theta}}(\boldsymbol{\theta},\boldsymbol{\theta})\leq\kappa^{2}\right\}.$ (12)
The CoRe describes all points for which the predictive covariance $s_{\boldsymbol{\Theta}}(\cdot,\cdot)$ is smaller or equal to a threshold $\kappa^{2}$, which is a hyperparameter of the model. For sufficiently small $\kappa$, an appropriate kernel function, and a weak prior, i.e., enough previously observed points, every point in the CoRe can be regarded as “already observed” because the uncertainty of the GP is small enough. Leveraging the concept of CoRe, Nicoli et al. [9] proposed the following acquisition function
$a^{\mathrm{EMICoRe}}\left(\boldsymbol{\Theta}^{\prime}\right)=\frac{1}{M}\left\langle\max\left(0,\min_{\boldsymbol{\theta}\in\mathcal{Z}_{\boldsymbol{\Theta}}}f(\boldsymbol{\theta})-\min_{\boldsymbol{\theta}\in\mathcal{Z}_{\widetilde{\boldsymbol{\Theta}}}}f(\boldsymbol{\theta})\right)\right\rangle_{p(f(\cdot)\mid\boldsymbol{\Theta},\boldsymbol{y})},$ (13)
where $\widetilde{\boldsymbol{\Theta}}=\left(\boldsymbol{\Theta},\boldsymbol{\Theta}^{\prime}\right)\in\mathcal{X}^{N+M}$ denotes the augmented training set with the new input points $\boldsymbol{\Theta}^{\prime}\in\mathcal{X}^{M}$, and the previously observed ones $\boldsymbol{\Theta}^{\prime}\in\mathcal{X}^{N}$.
The function $\min_{\boldsymbol{\theta}\in\mathcal{Z}_{\boldsymbol{\Theta}}}f(\boldsymbol{\theta})$ is the estimated minimum of the CoRe, identified by all previously observed points $\boldsymbol{\Theta}$. Such value is drawn from the GP’s posterior distribution and identifies the estimated minimum energy for any choice of circuit parameters previously observed. The other term in Eq. 13, i.e., $\min_{\boldsymbol{\theta}\in\mathcal{Z}_{\widetilde{\boldsymbol{\Theta}}}}f(\boldsymbol{\theta})$, is the estimated minimum over the augmented training set $\widetilde{\boldsymbol{\Theta}}$. The new input points $\boldsymbol{\Theta}^{\prime}\in\mathcal{X}^{M}$ are treated as if they were observed. In fact, when a point is in the CoRe, the GP’s uncertainty is smaller than a given threshold hence allowing to make the reasonable assumption that the GP is highly confident in a close range from that point. Therefore, one could implicitly assume this point as known, i.e., as if it has been observed (even though no measurement on the quantum computer has been performed).
3.2.1 EMICoRe
Since the NFT algorithm is not designed to deal with the statistical noise arising from measuring on a quantum computer, Nicoli et al. [9] proposed to combine NFT with GP regression and BO, thus merging all the tools previously introduced in Sections 2.2, 3.1 and 3.2.
The proposed algorithm builds on the structure of the NFT algorithm but improves it with two crucial modifications:
  1. 1.
    Learnable shifts: Instead of taking a fixed shift $\alpha$ to identify two new (equidistant) candidate points, one uses the EMICoRe acquisition function from Section 3.2, to identify the most informative points to measure next.
  2. 2.
    Re-use of previous measurements: For every iteration the estimation for the minimum on the one-dimensional subspace (using Least Square Regression) is performed using the posterior mean of the GP. This implies that the GP will give better estimates as more measurements are made on the quantum computer and used to train the surrogate model.
These modifications make the proposed algorithm very different from NFT, where the estimation of the minimum on the subspace solely relies on the previous best point and the fixed shifts for the new measurements. Both modifications aim at making EMICoRe a more flexible yet resilient algorithm with respect to the noise of the quantum computer. The EMICoRe algorithm starts with $T_{\textrm{NFT}}$ steps of the baseline NFT algorithm, necessary to collect sufficient measurements to instantiate an informative GP to make predictions. After this initial phase, the EMICoRe algorithm starts learning the shifts, for each step, by using the previous measurements and storing the new measurements for future predictions. For a detailed description of the EMICoRe algorithm and its subroutine, we refer to the original paper [9].
Figure 2 visualizes the algorithmic procedure of the full EMICoRe algorithm. The objective function $E^{\ast}(\boldsymbol{\theta})$ (true energy) is represented by the red lines, while the red shaded area indicates the standard deviation of an measurements on the quantum computer. The blue lines indicate the predictive mean of the GP $\mu_{\boldsymbol{\Theta}}(\cdot)$, while the shaded area denotes its predictive uncertainty $\sqrt{s_{\boldsymbol{\Theta}}(\cdot,\cdot)}$, i.e., one standard deviation. The figure should be read from top left to bottom right following the black arrow.
The top left plot shows the optimization in the direction $\theta_{d}$ for $d\in\{1,\ldots,D\}$ [^2] Note the direction to be optimized is chosen sequentially until $d=D$, after which the algorithm restarts from $d=0$.. The quantity ${\boldsymbol{\hat{\theta}}}^{(t)}$ denotes the set of parameters which give the lowest energy $E^{\ast}({\boldsymbol{\hat{\theta}}}^{(t)})$ after the $t-$th optimization step. Note that at this point the GP has access to a dataset of previous measurements $\{\boldsymbol{\Theta}^{(t)},\boldsymbol{\tilde{E}}^{(t)}\}$. This explains why the blue line, and the corresponding uncertainty, already give a good prediction of the red curve along direction $d$, even though no optimization has been performed yet.
The EMICoRe subroutine is now responsible for finding the promising points to be measured next, i.e., $\boldsymbol{\Theta}^{\prime}$ in Eq. 7 with the shifts $\alpha$ being different for $\theta_{1}^{\prime}$ and $\theta_{2}^{\prime}$. First, the subroutine divides the subspace into a grid of equally spaced points, as represented by the grid lines at the bottom of the second plot. Second, it prepares all possible combinations of candidate pairs $\boldsymbol{\Theta}^{\prime}_{k}=\{\boldsymbol{\theta}^{\prime}_{k,2},\boldsymbol{\theta}^{\prime}_{k,2}\}$. These points are added to the base-GP by pairs so that a discretized CoRe for each pair can be computed (orange vertical lines). The subroutine then computes the acquisition function for each candidate pair $\boldsymbol{\Theta}^{\prime}_{k}$ using the respective CoRe. The pair, which maximizes the acquisition function, see Eq. 13, is returned as $\boldsymbol{\Theta}^{\prime}$. According to the algorithm, this pair represents the points which are the most promising to be observed next. The fifth plot shows the measurements on the quantum computer at the suggested points, which are indicated by the blue crosses. The GP is finally updated using the new measurements— note the change of the blue curve and the shaded area in the plot. Finally, the GP’s posterior mean is computed for three equidistant points on the 1D-subspace. These points are then used to fit a cosine function according to the functional form of the subspace in Eq. 6. This fitted cosine is then used to obtain the analytic minimum of the 1D-subspace.
We recall that because of the special form of the VQE-kernel and the design of the EMICoRe algorithm, the functional sampled from the GP, i.e, the posterior mean, follows the same functional of Eq. 4. Therefore, fitting a sinusoidal function to the GP’s posterior is a valid assumption.
Figure 2: Visualization of the EMICoRe algorithm. We refer to the main text for more details.
4 Numerical Experiments
In the following, we perform some numerical experiments to demonstrate the performance of EMICoRe in the presence of hardware noise. Our experiments focus on the quantum Heisenberg Hamiltonian, a typical Hamiltonian for benchmarking the performance of VQEs [16]
$H=-\left[\sum_{j=1}^{Q-1}\left(J_{X}X_{j}X_{j+1}+J_{Y}Y_{j}Y_{j+1}+J_{Z}Z_{j}Z_{j+1}\right)+\sum_{j=1}^{Q}\left(h_{X}X_{j}+h_{Y}Y_{j}+h_{Z}Z_{j}\right)\right],$ (14)
where the index of the Pauli matrices $\{X_{j},Y_{j},Z_{j}\}$ refers to the qubit that the Pauli gate is applied to, and open boundary conditions are considered. In this work, we chose the couplings to be at the critical point with $(J_{X},J_{Y},J_{Z})=(-1,0,0)$ and $(h_{X},h_{Y},h_{Z})=(0,0,-1)$. The resulting Hamiltonian is also known as the Ising Hamiltonian at criticality.
In order to assess the performance of the VQE during the optimization, we mainly look at two metrics: energy and fidelity. The former is the lowest energy reached by the VQE during the optimization process, i.e., the expectation value of the Hamiltonian with respect to the variational state. The fidelity, measures the overlap between the variational (VQE) state and the true ground-state obtained with exact diagonalization. We refer to [9] for more details on the metrics we used. The ground-state energy $\underline{E}^{\ast}=\langle\psi_{\mathrm{GS}}|H|\psi_{\mathrm{GS}}\rangle$ together with the corresponding ground-state wave function $|\psi_{\mathrm{GS}}\rangle$ can be obtained in our case, because the size of the systems we tackle is sufficiently small. This is important as it gives us access to a ground truth for benchmarking different algorithms, i.e., NFT vs EMICoRe. To test the stability of the optimization with respect to initialization, we initialize several independent runs with different seeds. Specifically, we use $50$ runs for the experiments without Hardware noise and $10$ runs for the experiments with Hardware noise, see Section 4.1 for more details. In all plots we always report the median (50th percentile) as a solid line and the 25th and 75th percentiles being the lower and upper bounds respectively (shaded regions around the solid line). The results are plotted as a function of the total number of measurements performed on the quantum computer.
For all simulations, we use an $L$-layered and fully connected Efficient SU(2) circuit with $D=2(L+1)\cdot Q$ total angular parameters. The simulations of the quantum computer are performed with the Qiskit [17] library, while the Python implementation of the EMICoRe algorithm has been adapted from the original code available on GitHub [18].
All the parameters for EMICoRe are identical to the parameters chosen for the simulation of Fig. 7 in the original paper [9]. As noted in Footnote 1, we used $\pi/3$ as the shift for NFT. Moreover, in contrast to the original paper [7], we did not include a stabilization step, as our results showed that its use often result in poorer performance. We leave further investigation of this phenomenon to future work.
(a) Classical simulation of noisy quantum hardware without error mitigation.
(b) Classical simulation of noisy quantum hardware with TREX error mitigation.
(c) Classical simulation of noisy quantum hardware with ZNE error mitigation.
Figure 3: Energy (left) and Fidelity (right) for NFT sequential (green) and EMICoRe (red). We show results for the critical Ising Hamiltonian using $Q=5$ qubits and $L=3$ layers, for an Efficient SU(2) quantum circuit, and $N_{\textrm{shots}}=1024$ measurement shots. The top row shows runs with simulated hardware noise and no error mitigation, while in the second and third rows results for classical simulations of noisy hardware are shown, with TREX and ZNE error mitigation scheme respectively. The statistics, e.g., the median (solid line), $25^{\mathrm{th}}$ and $75^{\mathrm{th}}$ percentiles (colored regions surrounding the median), are obtained with $50$ and $10$ independent seeded runs for experiments in Fig. 3(a) and Fig. 3(b), respectively. A trial density showing the distribution of the runs is shown right next to each plot.
4.1 Hardware Noise Analysis of EMICoRe
Previous work [9] analyzed EMICoRe’s performance when measurement shot noise is present. It was observed, that EMICoRe outperforms NFT more significantly for higher levels of measurement shot noise, i.e., less shots. In this proceeding, we investigate the algorithm’s performance when accounting for hardware noise as well.
In the following, the setup is kept constant for all experiments with $5$-qubits and $3$-layers of the EfficientSU2 circuit and the Ising Hamiltonian with open boundary conditions. All simulations are performed for a total of $600$ measurements  [^3] Note that two measurements are performed for each step, i.e., $600$ measurements correspond to $300$ optimization steps. and $N=1024$ shots per measurement are used.
In Fig. 3(a) we run experiments with simulated hardware noise and no error mitigation technique while in Fig. 3(b) and Fig. 3(c) we use twisted readout error extinction (TREX) [19] and zero noise extrapolation (ZNE) [20, 21] as error mitigation strategies. The hardware noise is simulated with the Fake5QV1 backend. The results in Fig. 3(a) are obtained by aggregating $50$ independent runs, while Figs. 3(b) and 3(c) uses only $10$ runs, as the feature was deprecated before more data could be aggregated. Table 1 summarizes the results from Fig. 3, with mean and standard deviations computed over multiple seeded trials, for simulations with and without hardware noise.
When hardware noise is considered, we observe that EMICoRe significantly outperforms NFT, see Fig. 3(a), converging faster and more reliably than NFT, e.g., the trial density on the right-hand-side is more peaked around the median. We empirically observe that NFT struggles significantly more with hardware noise compared to EMICoRe. Notably, neither algorithm achieves the ground-state energy; however, EMICoRe attains a fidelity of nearly $80\%$. When using error mitigation, EMICoRe outperforms NFT in both energy minimization and fidelity, being more stable with respect to different initializations, and achieving both lower energies and higher fidelities.
Table 1: Mean and standard deviation of energy and fidelity calculated at the last optimization step, for all simulations shown in Fig. 3. For the energy lower is better, for the fidelity higher is better.
EMICoRe NFT
Energy $f^{\ast}({\boldsymbol{\hat{\theta}}})$ Fidelity Energy $f^{\ast}({\boldsymbol{\hat{\theta}}})$ Fidelity
noise $(-5.472\pm 0.119)$ $(0.782\pm 0.04)$ $(-4.972\pm 0.821)$ $(0.641\pm 0.214)$
noise, TREX $(-5.624\pm 0.122)$ $(0.854\pm 0.051)$ $(-5.173\pm 0.543)$ $(0.622\pm 0.245)$
noise, ZNE $(-5.611\pm 0.059)$ $(0.814\pm 0.037)$ $(-5.03\pm 0.841)$ $(0.653\pm 0.226)$
5 Conclusion
In these proceedings, we leveraged a novel machine-learning-enhanced optimization of VQE algorithms for finding the ground state of the quantum Ising Hamiltonian in the presence of simulated hardware noise. This algorithm, termed EMICoRe [9], builds on Gaussian Process Regression, for noise-resilient estimation of the energy, and Bayesian Optimization, to predict the new candidate points to measure at different step in the optimization.
Our numerical experiments validated the capability of EMICoRe to outperform the state-of-the-art baseline, the NFT algorithm [7], in the presence of simulated hardware noise. Even in the presence of simulated hardware noise, EMICoRe is able to converge to lower energies compared to NFT and is more robust with respect to different initializations. This robustness highlights the potential of the proposed method for future applications to more complicated quantum systems, and under the challenging conditions of real NISQ devices. As a next step, we aim to scale up EMICoRe for larger systems and ultimately evaluate its performance on real quantum hardware.
Acknowledgments
The authors thank Christopher J. Anders and Shinichi Nakajima for insightful discussions. This project was supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) as part of the CRC 1639 NuMeriQS – project no. 511713970.

### v2

@@
 VQEs can approximate the ground state of quantum Hamiltonians by classical variational optimization of a parametrized quantum circuit.
-The quantum computer measures the energy of the quantum system for a given set of parameters, i.e., the objective of the minimization problem, and a classical optimization routine is used to find a better set of parameters.
+The quantum computer measures the energy of the quantum system for a given set of parameters, i.e., the objective of the minimization problem, while a classical optimization routine is used to find a better set of parameters.
 Efficient exploration of the parameter space is crucial for successful optimization.
@@
 In a recent work [9], Nicoli et al.
-proposed an optimization algorithm called NFT-with-EMICoRe (EMICoRe for short), which extends the NFT framework by incorporating Gaussian Process Regression (GPR) and Bayesian Optimization (BO) for a data-driven and potentially more noise-resilient sequential minimal optimization scheme.
+proposed an optimization algorithm called NFT-with-EMICoRe (EMICoRe for short), which extends the NFT framework by incorporating Gaussian Process Regression (GPR) and Bayesian Optimization (BO) for a data-driven and potentially more noise-resilient SMO scheme.
 However, this work [9] focused solely on measurement shot noise, neglecting the impact of quantum hardware noise, such as decoherence, depolarization, and crosstalk errors.
-In these proceedings, we build on prior works [7, 9] and demonstrate that EMICoRe outperforms the NFT baseline, even in the presence of hardware noise.
+In these proceedings, we build on prior works [7, 9] and demonstrate that EMICoRe outperforms the NFT baseline in the presence of simulated quantum hardware noise.
 We start in Sec.
@@
 }$ (3)
-The energy $E^{\ast}(\boldsymbol{\theta})$ is the true energy that that would be obtained if the measurement on the quantum computer were error-free.
-Typically, a quantum computer experiences both Gaussian noise, arising from the statistical nature of repeated measurements (i.e., measurement shot noise), and hardware noise, which results from imperfections in qubits, gates, and measurements [10].
+The energy $E^{\ast}(\boldsymbol{\theta})$ is the true energy that would be obtained if the quantum computer were noise-free.
+Typically, a quantum computer is subject to both Gaussian noise, arising from the statistical nature of repeated measurements (i.e., measurement shot noise), and hardware noise, which stems from imperfections in qubits, gates, and measurements [10].
 Both types of noise are encoded in the noise term $\varepsilon$ in Eq.
@@
 where $\boldsymbol{b}$ is a set of arbitrary coefficients and $\theta_{d}$ is a $D$-dimensional vector of angular parameters for the variational quantum circuit.
-In particular, this functional form allows the following observation: when keeping all but one parameter constant, i.e.,
+In particular, this functional form leads to the following observation: when keeping all but one parameter constant, i.e.,
 $\boldsymbol{\theta}^{(t)}_{d}=(\theta_{1},\ldots,\theta_{d-1},\theta_{d},\theta_{d+1},\ldots,\theta_{D})^{\top}\quad\mathrm{with}\quad\theta_{i\neq d}=\textrm{const.,}$., (5)
-the form of the objective function for this one-dimensional subspace of the whole optimization problem reduces to
+the form of the objective function for this one-dimensional subspace of the whole optimization problem reduces to [7]
 $E^{\ast(t)}(\boldsymbol{\theta}^{(t)}_{d})$ $=\sum_{k=1}^{K}h_{k}\left\langle\psi_{k}\left|U_{d}^{\dagger}\left(\theta^{(t)}_{d}\right)P_{k}U_{d}\left(\theta^{(t)}_{d}\right)\right|\psi_{k}\right\rangle$ (6)
@@
 where $a^{(t)}_{\ell,d}(\ell=1,2,3)$ denote parameters that are independent of $\theta^{(t)}_{d}$, while the superscript $(t)$ refers to the optimization step.
-As a result, it is enough to measure three points in the one-dimensional space to fit a function over the entire subspace and analytically determine the minimum.
+As a result, measuring three points in the one-dimensional subspace is sufficient to fit a function across the entire domain and analytically determine its minimum.
 The NFT optimization algorithm [7] uses this property and sequentially optimizes each parameter in the set $\boldsymbol{\theta}$ by requiring (at least) two measurements for each optimization step.
@@
 $\boldsymbol{\Theta}^{\prime}=\left(\boldsymbol{\theta}_{1}^{\prime},\boldsymbol{\theta}_{2}^{\prime}\right)=\left\{{\boldsymbol{\hat{\theta}}}^{(t-1)}-\alpha\boldsymbol{e}_{d},{\boldsymbol{\hat{\theta}}}^{(t-1)}+\alpha\boldsymbol{e}_{d}\right\}\,,$ (7)
-where $\boldsymbol{\hat{\theta}}^{(t-1)}$ denotes the previous best point from step $(t-1)$ [^1] The original paper from Nakanishi et al.
+where $\boldsymbol{\hat{\theta}}^{(t-1)}$ denotes the previous best point from step $(t-1)$.
+[^1] The original paper from Nakanishi et al.
 [7] proposed a shift $\alpha$ of $\frac{\pi}{2}$ for Eq.
@@
 In the work by Nicoli et al.
-[9] it was suggested, that equidistant points for probing the subspace should be used, i.e., a shift of $\frac{2\pi}{3}$ should be preferred.
+[9] it was suggested that equidistant points for probing the subspace should be used, i.e., a shift of $\frac{2\pi}{3}$ should be preferred.
 Later, it was derived by Anders et al.
@@
 In a private communication with the authors, it was reported that, because of a bug, the actual value used in the experiments of [9] was $\frac{\pi}{3}$ instead of $\frac{2\pi}{3}$ as reported in the paper.
-Although this did not lead to significant differences in the numerical experiments, in order to be consistent with the prior work, we will stick to the convention of using $\frac{\pi}{3}$ for NFT..
+Although this did not lead to significant differences in the numerical experiments, in order to be consistent with the prior work, we will stick to the convention of using $\frac{\pi}{3}$ for NFT.
 This approach is advantageous because it enables analytical optimization of each parameter within its respective one-dimensional subspace, eliminating the need to compute gradients as required by gradient-based methods, such as SPSA [12, 13].
@@
 2.2 Gaussian Processes and Bayesian Optimization
-The NFT algorithm can find the exact analytical minimum within the one-dimensional subspace if and only the measurements on the quantum computer are exact, i.e., an infinite number of shots and no hardware noise.
-However, this is not a real-world scenario.
-In fact, due to the unavoidable hardware noise on current NISQ hardware, and the inherent statistical uncertainty from repeated measurements (shots), it may be beneficial to enhance existing algorithms with frameworks like Bayesian Optimization (BO) [14] and Gaussian Process Regression (GPR) [15].
+The NFT algorithm can find the exact analytical minimum within the one-dimensional subspace if and only if the state preparation and measurements on the quantum computer are exact—that is, if an infinite number of shots are taken and no hardware noise is present.
+In a real-world scenario, due to the unavoidable hardware noise on current NISQ hardware and the inherent statistical uncertainty from repeated measurements (shots), it may be beneficial to enhance existing algorithms with frameworks like Bayesian Optimization (BO) [14] and Gaussian Process Regression (GPR) [15].
 These techniques, widely established in machine learning, are often well-suited to deal with noisy data [15].
-Let $E^{\ast}(\cdot):\mathcal{X}\mapsto\mathbb{R}$ be an unknown (black-box) objective function to be minimized (i.e., the energy of a VQE in our case).
+Let $E^{\ast}(\cdot):\mathcal{X}\mapsto\mathbb{R}$ be an unknown (black-box) objective function to be minimized (i.e., the energy in our case).
 With the aid of a function that suggests promising new points to observe, often referred to as acquisition function [14], one iteratively approximates the target objective with a surrogate function $E(\boldsymbol{\theta})\approx E^{\ast}(\boldsymbol{\theta})$, sampled from the GP.
 At each optimization step, after new points are measured, those are added to the previous set of measured points, used to update the GP.
-A common choice as a surrogate model is a GP regression model with a one-dimensional Gaussian likelihood and GP prior [9]
+A common choice for a surrogate model is a GP regression model with a one-dimensional Gaussian likelihood and a GP prior [9],
 $p(\tilde{E}\mid\boldsymbol{\theta},E(\cdot))=\mathcal{N}_{1}\left(\tilde{E};E(\boldsymbol{\theta}),\sigma^{2}\right),\quad p(E(\cdot))=\operatorname{GP}(E(\cdot);\nu(\cdot),k(\cdot,\cdot)).$ (8)
-Here, $\sigma^{2}$ is the variance of the measurement noise $\varepsilon$, and the prior mean function $\nu(\cdot)$ is set to 00 in this work.
-In the case of VQEs a zero-mean prior is a reasonable choice because the target function $E^{\ast}(\boldsymbol{\theta})$ one aims to model is a linear combination of functions centered around zero (up to a potential constant offset which stems from Eq.
+Here, $\sigma^{2}$ is the variance of the noise $\varepsilon$ defined in Eq.
+2, and the prior mean function $\nu(\cdot)$ is set to 00 in this work.
+In the case of VQEs, a zero-mean prior is a reasonable choice because the target function $E^{\ast}(\boldsymbol{\theta})$, which one aims to model, is a linear combination of functions centered around zero (up to a potential constant offset stemming from Eq.
 4).
 The function $k(\cdot,\cdot)$ is the kernel (covariance) function encoding the expected correlation between two measured points.
-We refer to  Section 3.1 for more details on the kernel function we use in this proceeding.
+We refer to  Section 3.1 for more details on the kernel function we use in these proceedings.
 The GP is trained on noisy measurements $\tilde{E}=E^{\ast}(\boldsymbol{\theta})+\varepsilon$ that were collected throughout previous optimization steps.
@@
 Thus, in Ref.
-[9] it was shown that for arbitrary $M$ test points $\boldsymbol{\Theta}^{\prime}=\left(\boldsymbol{\theta}_{1}^{\prime},\ldots,\boldsymbol{\theta}^{\prime}{}_{M}\right)\in\mathcal{X}^{M}$, the posterior, or predictive distribution of the function values $\boldsymbol{E}^{\prime}=\left(E\left(\boldsymbol{\theta}_{1}^{\prime}\right),\ldots,E\left(\boldsymbol{\theta}_{M}^{\prime}\right)\right)^{\top}\in\mathbb{R}^{M}$ is the $M$-dimensional Gaussian with its mean and covariance analytically given
-$\begin{gathered}p\left(\boldsymbol{E}^{\prime}\mid\boldsymbol{\Theta},\boldsymbol{\tilde{E}}\right)=\mathcal{N}_{M}\left(\boldsymbol{E}^{\prime};\boldsymbol{\mu}_{\boldsymbol{\Theta}}^{\prime},\boldsymbol{S}_{\boldsymbol{\Theta}}^{\prime}\right),\quad\text{ where }\\ \boldsymbol{\mu}_{\boldsymbol{\Theta}}^{\prime}=\boldsymbol{K}^{\prime\top}\left(\boldsymbol{K}+\sigma^{2}\boldsymbol{I}_{N}\right)^{-1}\boldsymbol{\tilde{E}},\quad\boldsymbol{S}_{\boldsymbol{\Theta}}^{\prime}=\boldsymbol{K}^{\prime\prime}-\boldsymbol{K}^{\prime\top}\left(\boldsymbol{K}+\sigma^{2}\boldsymbol{I}_{N}\right)^{-1}\boldsymbol{K}^{\prime}.\end{gathered}$ (9)
-The notation with a prime refers to new, unseen, points with the corresponding energies predicted by the GP.
-Here, $\boldsymbol{K}=k(\boldsymbol{\Theta},\boldsymbol{\Theta})\in\mathbb{R}^{N\times N},\boldsymbol{K}^{\prime}=k\left(\boldsymbol{\Theta},\boldsymbol{\Theta}^{\prime}\right)\in\mathbb{R}^{N\times M}$, and $\boldsymbol{K}^{\prime\prime}=k\left(\boldsymbol{\Theta}^{\prime},\boldsymbol{\Theta}^{\prime}\right)\in\mathbb{R}^{M\times M}$ are the train, train-test, and test kernel matrices, respectively, where $k\left(\boldsymbol{\Theta},\boldsymbol{\Theta}^{\prime}\right)$ denotes the kernel matrix evaluated at each column of $\boldsymbol{\Theta}$ and $\boldsymbol{\Theta}^{\prime}$ such that $\left(k\left(\boldsymbol{\Theta},\boldsymbol{\Theta}^{\prime}\right)\right)_{n,m}=k\left(\boldsymbol{\theta}_{n},\boldsymbol{\theta}_{m}\right)$.
+[9], it was shown that for arbitrary $M$ test points $\boldsymbol{\Theta}^{\prime}=\left(\boldsymbol{\theta}_{1}^{\prime},\ldots,\boldsymbol{\theta}^{\prime}{}_{M}\right)\in\mathcal{X}^{M}$, the posterior, or predictive distribution of the function values $\boldsymbol{E}^{\prime}=\left(E\left(\boldsymbol{\theta}_{1}^{\prime}\right),\ldots,E\left(\boldsymbol{\theta}_{M}^{\prime}\right)\right)^{\top}\in\mathbb{R}^{M}$, is an $M$-dimensional Gaussian,
+$\begin{gathered}p\left(\boldsymbol{E}^{\prime}\mid\boldsymbol{\Theta},\boldsymbol{\tilde{E}}\right)=\mathcal{N}_{M}\left(\boldsymbol{E}^{\prime};\boldsymbol{\mu}_{\boldsymbol{\Theta}}^{\prime},\boldsymbol{S}_{\boldsymbol{\Theta}}^{\prime}\right),\end{gathered}$ (9)
+with mean and covariance analytically given by
+$\begin{gathered}\boldsymbol{\mu}_{\boldsymbol{\Theta}}^{\prime}=\boldsymbol{K}^{\prime\top}\left(\boldsymbol{K}+\sigma^{2}\boldsymbol{I}_{N}\right)^{-1}\boldsymbol{\tilde{E}},\quad\boldsymbol{S}_{\boldsymbol{\Theta}}^{\prime}=\boldsymbol{K}^{\prime\prime}-\boldsymbol{K}^{\prime\top}\left(\boldsymbol{K}+\sigma^{2}\boldsymbol{I}_{N}\right)^{-1}\boldsymbol{K}^{\prime}.\end{gathered}$ (10)
+The notation with a prime refers to new, unseen points with the corresponding energies predicted by the GP.
+Here, $\boldsymbol{K}=k(\boldsymbol{\Theta},\boldsymbol{\Theta})\in\mathbb{R}^{N\times N},\boldsymbol{K}^{\prime}=k\left(\boldsymbol{\Theta},\boldsymbol{\Theta}^{\prime}\right)\in\mathbb{R}^{N\times M}$, and $\boldsymbol{K}^{\prime\prime}=k\left(\boldsymbol{\Theta}^{\prime},\boldsymbol{\Theta}^{\prime}\right)\in\mathbb{R}^{M\times M}$ are the train, train-test, and test kernel matrices, respectively, where $k\left(\boldsymbol{\Theta},\boldsymbol{\Theta}^{\prime}\right)$ denotes the kernel matrix evaluated at each column of $\boldsymbol{\Theta}$ and $\boldsymbol{\Theta}^{\prime}$, such that $\left(k\left(\boldsymbol{\Theta},\boldsymbol{\Theta}^{\prime}\right)\right)_{n,m}=k\left(\boldsymbol{\theta}_{n},\boldsymbol{\theta^{\prime}}_{m}\right)$.
 Moreover, $\boldsymbol{I}_{N}\in\mathbb{R}^{N\times N}$ denotes the identity matrix.
@@
 9 denotes which input points were used to train the GP.
-Note that the predictive mean and covariance are given in their vector or matrix form, respectively.
-The predictive mean can be computed as a function of one arbitrary input vector $\mu_{\boldsymbol{\Theta}}(\boldsymbol{\theta}^{\prime})=\boldsymbol{K}^{\prime\top}\left(\boldsymbol{K}+\sigma^{2}\boldsymbol{I}_{N}\right)^{-1}\boldsymbol{\tilde{E}}$ with $\boldsymbol{K}^{\prime}=k\left(\boldsymbol{\Theta},\boldsymbol{\theta}^{\prime}\right)\in\mathbb{R}^{N}$.
-The computation of a single predictive covariance $s_{\boldsymbol{\Theta}}(\cdot,\cdot)$ follows analogously.
-For more details on how to compute the predictive mean and variance we refer to Ref.
+Note that the predictive mean and covariance are given in their vector and matrix form, respectively.
+The predictive mean can be computed as a function of one arbitrary test point $\boldsymbol{\theta}^{\prime}$, i.e., $\mu_{\boldsymbol{\Theta}}(\boldsymbol{\theta}^{\prime})=\boldsymbol{K}^{\prime\top}\left(\boldsymbol{K}+\sigma^{2}\boldsymbol{I}_{N}\right)^{-1}\boldsymbol{\tilde{E}}$ with $\boldsymbol{K}^{\prime}=k\left(\boldsymbol{\Theta},\boldsymbol{\theta}^{\prime}\right)\in\mathbb{R}^{N}$.
+The computation of a single predictive covariance $s_{\boldsymbol{\Theta}}(\cdot,\cdot)$ with $\boldsymbol{S}_{\boldsymbol{\Theta}}^{\prime}=s_{\boldsymbol{\Theta}}\left(\boldsymbol{\Theta}^{\prime},\boldsymbol{\Theta}^{\prime}\right)\in\mathbb{R}^{M\times M}$ follows analogously.
+For more details on how to compute the predictive mean and variance, we refer to Ref.
 [9].
-The knowledge gathered by the GP during optimization, as the number of measurements increases, can not only be exploited for regression but can also be leveraged to choose the next points to observe.
-This is achieved by combining GPR and BO and solving the following maximization problem
-$\max_{\boldsymbol{\Theta^{\prime}}}{a_{\boldsymbol{\Theta}^{(t-1)}}(\boldsymbol{\Theta^{\prime}})}\,$
+As the number of measurements increases, the knowledge gained by the GP during optimization can be used not only for regression but also to choose the next points to observe.
+This is achieved by combining GPR and BO and solving the following maximization problem:
+$\max_{\boldsymbol{\Theta^{\prime}}}{a_{\boldsymbol{\Theta}^{(t-1)}}(\boldsymbol{\Theta^{\prime}})}\,,$
 where $a_{\boldsymbol{\Theta}^{(t-1)}}(\cdot)$ is the so-called acquisition function.
-This latter uses a GP model trained on previously observed data pairs $\{\boldsymbol{\Theta}^{(t-1)},\boldsymbol{\tilde{E}}\}$, collected over the prior $(t-1)$ iterations, to assess how promising a set of new input points $\boldsymbol{\Theta^{\prime}}$ may be.
-Points that maximize the acquisition function are likely to be the most informative points to measure at the next iteration, such that the energy is minimized.
+This function uses a GP model trained on previously observed data pairs $\{\boldsymbol{\Theta}^{(t-1)},\boldsymbol{\tilde{E}}\}$, collected over the prior $(t-1)$ iterations, to assess how promising a set of new input points $\boldsymbol{\Theta^{\prime}}$ may be.
+The points that maximize the acquisition function are likely to be the most informative for the next iteration.
 3 Method
-In this section, we leverage the functional form of the VQE objective as derived in [7] and we construct a physics-informed kernel suited for GPR.
-Furthermore, we recall an innovative type of acquisition function and the concept of confident region (CoRe) as introduced in [9].
-Finally, in Section 3.2.1, we show how the NFT baseline algorithm can be improved to achieve a more flexible and noise-resilient algorithm when combined with BO and GPR.
+In Section 3.1, we leverage the functional form of the VQE objective [7] to construct a physics-informed kernel suited for GPR [9].
+In Section 3.2, we introduce a novel type of acquisition function called Expected Maximum Improvement over Confident Regions (EMICoRe) [9].
+In Section 3.2.1, we combine these concepts and present the NFT-with-EMICoRe algorithm, called EMICoRe algorithm for short [9].
+This algorithm demonstrates how the NFT baseline algorithm can be improved to achieve a more flexible and noise-resilient algorithm when combined with BO and GPR.
 3.1 VQE Kernel
 A GP needs to be equipped with a kernel (covariance) function.
-While there are many types of such functions, one popular example is the Gaussian-RBF kernel [15], which encodes into the GP the belief that input vectors that are close in the input space strongly correlate compared to points that are far apart, with the correlation strength decaying exponentially with the distance.
-Although this might be a reasonable assumption, an exponential drop in correlation, like in the Gaussian-RBF kernel, may not be representative of the underlying features of every problem.
-For this reason, wisely choosing the kernel for a GP regression task can substantially affect the performance.
+While there are many types of such functions, one popular example is the Gaussian-RBF kernel [15].
+This kernel encodes the assumption that input vectors that are close in the input space exhibit stronger correlations compared to those that are distant, with the correlation strength decaying exponentially as a function of the distance.
+While this may be a reasonable assumption, an exponential decay in correlation, as seen in the Gaussian-RBF kernel, may not accurately reflect the underlying features of every problem.
+For this reason, the choice of kernel in a GPR task can significantly affect performance.
 In Ref.
@@
 Both $\sigma_{0}$ and $\gamma^{2}$ are hyperparameters of the kernel.
-It was shown that the VQE kernel can be decomposed as
-$k^{\mathrm{VQE}}\left(\boldsymbol{\theta},\boldsymbol{\theta}^{\prime}\right)=\boldsymbol{\phi}(\boldsymbol{\theta})^{\top}\boldsymbol{\phi}(\boldsymbol{\theta}^{\prime}),\quad\textrm{with}\quad\boldsymbol{\phi}(\boldsymbol{\theta})=\sigma_{0}\left(1+\gamma^{2}\right)^{-D/2}\left[\otimes_{d=1}^{D}\left(\gamma,\cos\theta_{d},\sin\theta_{d}\right)^{\top}\right]$ (11)
-where the feature maps $\boldsymbol{\phi}(\boldsymbol{\theta})$ fulfill the same VQE functional form derived in Eq.
+In [9], it was shown that the VQE kernel can be decomposed as
+$k^{\mathrm{VQE}}\left(\boldsymbol{\theta},\boldsymbol{\theta}^{\prime}\right)=\boldsymbol{\phi}(\boldsymbol{\theta})^{\top}\boldsymbol{\phi}(\boldsymbol{\theta}^{\prime}),\quad\textrm{with}\quad\boldsymbol{\phi}(\boldsymbol{\theta})=\sigma_{0}\left(1+\gamma^{2}\right)^{-D/2}\left[\otimes_{d=1}^{D}\left(\gamma,\cos\theta_{d},\sin\theta_{d}\right)^{\top}\right],$ (12)
+where the feature maps $\boldsymbol{\phi}(\boldsymbol{\theta})$ fulfill the same VQE functional form as derived in Eq.
 4.
-Since a GP regression model is only defined in terms of inner products of the feature maps, one can replace all of these occurrences with the kernel described above – a procedure often known as kernel trick [15].
-Using the VQE kernel function instead of, e.g., the RBF-kernel, ensures that the GP samples functions that match the required functional form from Eq.
+Since a GP regression model is only defined in terms of inner products of the feature maps, one just requires access to the kernel form without explicitly specifying $\boldsymbol{\phi}$ —a procedure known as kernel trick [15].
+By using the VQE kernel instead of, for example, the Gaussian-RBF kernel, the GP is ensured to sample functions that align with the required functional form from Eq.
 4.
 Consequently, as described in Eq.
-6, when optimizing on a one-dimensional subspace, i.e., changing only one parameter at a time, three (noiseless) observed points allow to uniquely identify the target function [7] over the entire one-dimensional subspace.
-However, in the presence of statistical uncertainty due to repeated measurements (shots), the target function in the subspace can be identified up to some statistical uncertainty.
-Therefore, the GP will be able to provide energy estimates, and their corresponding uncertainties, for new measurements in the subspace.
+6, when optimizing over a one-dimensional subspace—i.e., varying only one parameter at a time—three noiseless observed points are sufficient to uniquely identify the target function [7] across the entire subspace.
+However, in the presence of measurement shot noise, the target function in the subspace can only be identified up to some statistical uncertainty.
+Therefore, the GP provides not only energy estimates but also their corresponding uncertainties for new measurements in the subspace.
 Furthermore, it has been shown that the decomposition of the VQE kernel in Eq.
 12 can be generalized so that each entry of $\boldsymbol{\theta}$ may parametrize multiple gates at once [9].
-This property becomes crucial if the Hamiltonian is invariant under certain symmetries, which results in multiple parameters being identical and requiring simultaneous rotation.
-3.2 EMICoRe: Expected Maximum Improvement over Confident Regions
-To perform Bayesian Optimization (BO) in combination with GPR, one should pick a suitable acquisition function to identify the most promising points to measure at each optimization step.
+This property becomes crucial if the Hamiltonian is invariant under specific symmetries, leading to multiple parameters being identical and requiring simultaneous updates.
+3.2 Expected Maximum Improvement over Confident Regions
+To perform BO in combination with GPR, one should pick a suitable acquisition function to identify the most promising points to measure at each optimization step.
 A novel acquisition function termed Expected Maximum Improvement over Confident Regions (EMICoRe) was introduced in the context of GPR for VQEs [9].
@@
 $\mathcal{Z}_{\boldsymbol{\Theta}}=\left\{\boldsymbol{\theta}\in\mathcal{X};s_{\boldsymbol{\Theta}}(\boldsymbol{\theta},\boldsymbol{\theta})\leq\kappa^{2}\right\}.$ (13)
-The CoRe describes all points for which the predictive covariance $s_{\boldsymbol{\Theta}}(\cdot,\cdot)$ is smaller or equal to a threshold $\kappa^{2}$, which is a hyperparameter of the model.
-For sufficiently small $\kappa$, an appropriate kernel function, and a weak prior, i.e., enough previously observed points, every point in the CoRe can be regarded as “already observed” because the uncertainty of the GP is small enough.
+The CoRe describes all points for which the predictive covariance $s_{\boldsymbol{\Theta}}(\cdot,\cdot)$ is smaller than or equal to a threshold $\kappa^{2}$, which is a hyperparameter of the model.
+For sufficiently small $\kappa$, an appropriate kernel function, and a weak prior—i.e., enough previously observed points—every point in the CoRe can be regarded as “already observed” due to the sufficiently small uncertainty of the GP.
 Leveraging the concept of CoRe, Nicoli et al.
-[9] proposed the following acquisition function
-$a^{\mathrm{EMICoRe}}\left(\boldsymbol{\Theta}^{\prime}\right)=\frac{1}{M}\left\langle\max\left(0,\min_{\boldsymbol{\theta}\in\mathcal{Z}_{\boldsymbol{\Theta}}}f(\boldsymbol{\theta})-\min_{\boldsymbol{\theta}\in\mathcal{Z}_{\widetilde{\boldsymbol{\Theta}}}}f(\boldsymbol{\theta})\right)\right\rangle_{p(f(\cdot)\mid\boldsymbol{\Theta},\boldsymbol{y})},$ (13)
-where $\widetilde{\boldsymbol{\Theta}}=\left(\boldsymbol{\Theta},\boldsymbol{\Theta}^{\prime}\right)\in\mathcal{X}^{N+M}$ denotes the augmented training set with the new input points $\boldsymbol{\Theta}^{\prime}\in\mathcal{X}^{M}$, and the previously observed ones $\boldsymbol{\Theta}^{\prime}\in\mathcal{X}^{N}$.
-The function $\min_{\boldsymbol{\theta}\in\mathcal{Z}_{\boldsymbol{\Theta}}}f(\boldsymbol{\theta})$ is the estimated minimum of the CoRe, identified by all previously observed points $\boldsymbol{\Theta}$.
-Such value is drawn from the GP’s posterior distribution and identifies the estimated minimum energy for any choice of circuit parameters previously observed.
-The other term in Eq.
-13, i.e., $\min_{\boldsymbol{\theta}\in\mathcal{Z}_{\widetilde{\boldsymbol{\Theta}}}}f(\boldsymbol{\theta})$, is the estimated minimum over the augmented training set $\widetilde{\boldsymbol{\Theta}}$.
-The new input points $\boldsymbol{\Theta}^{\prime}\in\mathcal{X}^{M}$ are treated as if they were observed.
-In fact, when a point is in the CoRe, the GP’s uncertainty is smaller than a given threshold hence allowing to make the reasonable assumption that the GP is highly confident in a close range from that point.
-Therefore, one could implicitly assume this point as known, i.e., as if it has been observed (even though no measurement on the quantum computer has been performed).
-3.2.1 EMICoRe
-Since the NFT algorithm is not designed to deal with the statistical noise arising from measuring on a quantum computer, Nicoli et al.
+[9] proposed the following acquisition function:
+$a^{\mathrm{EMICoRe}}\left(\boldsymbol{\Theta}^{\prime}\right)=\frac{1}{M}\left\langle\max\left(0,\min_{\boldsymbol{\theta}\in\mathcal{Z}_{\boldsymbol{\Theta}}}E(\boldsymbol{\theta})-\min_{\boldsymbol{\theta}\in\mathcal{Z}_{\widetilde{\boldsymbol{\Theta}}}}E(\boldsymbol{\theta})\right)\right\rangle_{p(E(\cdot)\mid\boldsymbol{\Theta},\boldsymbol{\tilde{E}})},$ (14)
+where $\widetilde{\boldsymbol{\Theta}}=\left(\boldsymbol{\Theta},\boldsymbol{\Theta}^{\prime}\right)\in\mathcal{X}^{N+M}$ denotes the augmented training set with the new input points $\boldsymbol{\Theta}^{\prime}\in\mathcal{X}^{M}$ and the previously observed ones $\boldsymbol{\Theta}\in\mathcal{X}^{N}$.
+[…]
+3.2.1 The EMICoRe algorithm
+To address the challenges posed by noisy measurements on quantum hardware, Nicoli et al.
 [9] proposed to combine NFT with GP regression and BO, thus merging all the tools previously introduced in Sections 2.2, 3.1 and 3.2.
-The proposed algorithm builds on the structure of the NFT algorithm but improves it with two crucial modifications:
+The proposed EMICoRe algorithm builds upon the structure of the NFT algorithm, incorporating two key modifications:
   1. 1.
-    Learnable shifts: Instead of taking a fixed shift $\alpha$ to identify two new (equidistant) candidate points, one uses the EMICoRe acquisition function from Section 3.2, to identify the most informative points to measure next.
+    Learnable shifts: Instead of taking a fixed shift $\alpha$ to identify two new (equidistant) candidate points, see Eq.
+7, the EMICoRe algorithm uses the EMICoRe acquisition function from Section 3.2 to identify the most informative points to measure next.
   2. 2.
-    Re-use of previous measurements: For every iteration the estimation for the minimum on the one-dimensional subspace (using Least Square Regression) is performed using the posterior mean of the GP.
-This implies that the GP will give better estimates as more measurements are made on the quantum computer and used to train the surrogate model.
-These modifications make the proposed algorithm very different from NFT, where the estimation of the minimum on the subspace solely relies on the previous best point and the fixed shifts for the new measurements.
-Both modifications aim at making EMICoRe a more flexible yet resilient algorithm with respect to the noise of the quantum computer.
-The EMICoRe algorithm starts with $T_{\textrm{NFT}}$ steps of the baseline NFT algorithm, necessary to collect sufficient measurements to instantiate an informative GP to make predictions.
-After this initial phase, the EMICoRe algorithm starts learning the shifts, for each step, by using the previous measurements and storing the new measurements for future predictions.
+    Re-use of previous measurements: At each iteration, the minimum on the one-dimensional subspace is estimated using Least Square Regression with the posterior mean of the GP.
+This implies that as more measurements are made on the quantum computer and incorporated into the surrogate model, the GP will provide increasingly accurate estimates.
+These modifications distinguish the proposed algorithm from NFT, where the minimum estimation on the subspace depends solely on the previous best point and the fixed shifts for the new measurements.
+Both modifications aim to make EMICoRe a more flexible and resilient algorithm in the presence of quantum noise.
+The EMICoRe algorithm starts with $T_{\textrm{NFT}}$ steps of the baseline NFT algorithm to gather sufficient measurements for instantiating an informative GP that can make predictions.
+After this initial phase, the EMICoRe algorithm begins learning the shifts at each step, using the previous measurements and storing the new measurements for future predictions.
 For a detailed description of the EMICoRe algorithm and its subroutine, we refer to the original paper [9].
 Figure 2 visualizes the algorithmic procedure of the full EMICoRe algorithm.
-The objective function $E^{\ast}(\boldsymbol{\theta})$ (true energy) is represented by the red lines, while the red shaded area indicates the standard deviation of an measurements on the quantum computer.
+The objective function $E^{\ast}(\boldsymbol{\theta})$ (true energy) is represented by the red lines, while the red shaded area indicates the standard deviation of a measurement on the quantum computer.
 The blue lines indicate the predictive mean of the GP $\mu_{\boldsymbol{\Theta}}(\cdot)$, while the shaded area denotes its predictive uncertainty $\sqrt{s_{\boldsymbol{\Theta}}(\cdot,\cdot)}$, i.e., one standard deviation.
@@
 The top left plot shows the optimization in the direction $\theta_{d}$ for $d\in\{1,\ldots,D\}$ [^2] Note the direction to be optimized is chosen sequentially until $d=D$, after which the algorithm restarts from $d=0$..
-The quantity ${\boldsymbol{\hat{\theta}}}^{(t)}$ denotes the set of parameters which give the lowest energy $E^{\ast}({\boldsymbol{\hat{\theta}}}^{(t)})$ after the $t-$th optimization step.
-Note that at this point the GP has access to a dataset of previous measurements $\{\boldsymbol{\Theta}^{(t)},\boldsymbol{\tilde{E}}^{(t)}\}$.
-This explains why the blue line, and the corresponding uncertainty, already give a good prediction of the red curve along direction $d$, even though no optimization has been performed yet.
+The quantity ${\boldsymbol{\hat{\theta}}}^{(t)}$ denotes the set of parameters which give the lowest energy $E({\boldsymbol{\hat{\theta}}}^{(t)})$ from the GP after the $t$-th optimization step.
+Note that at this point, the GP has access to a dataset of previous measurements $\{\boldsymbol{\Theta}^{(t)},\boldsymbol{\tilde{E}}^{(t)}\}$.
+This explains why the blue line, along with the corresponding uncertainty, provides a good prediction of the red curve along direction $d$, even before any optimization has been performed.
 The EMICoRe subroutine is now responsible for finding the promising points to be measured next, i.e., $\boldsymbol{\Theta}^{\prime}$ in Eq.
@@
 First, the subroutine divides the subspace into a grid of equally spaced points, as represented by the grid lines at the bottom of the second plot.
-Second, it prepares all possible combinations of candidate pairs $\boldsymbol{\Theta}^{\prime}_{k}=\{\boldsymbol{\theta}^{\prime}_{k,2},\boldsymbol{\theta}^{\prime}_{k,2}\}$.
-These points are added to the base-GP by pairs so that a discretized CoRe for each pair can be computed (orange vertical lines).
+Second, the subroutine prepares all possible combinations of candidate pairs $\boldsymbol{\Theta}^{\prime}_{k}=\{\boldsymbol{\theta}^{\prime}_{k,2},\boldsymbol{\theta}^{\prime}_{k,2}\}$.
+These points are added to the base-GP in pairs, allowing a discretized CoRe to be computed for each pair (orange vertical lines).
 The subroutine then computes the acquisition function for each candidate pair $\boldsymbol{\Theta}^{\prime}_{k}$ using the respective CoRe.
-The pair, which maximizes the acquisition function, see Eq.
-13, is returned as $\boldsymbol{\Theta}^{\prime}$.
-According to the algorithm, this pair represents the points which are the most promising to be observed next.
-The fifth plot shows the measurements on the quantum computer at the suggested points, which are indicated by the blue crosses.
+The pair that maximizes the acquisition function, as described in Eq.
+14, is returned as $\boldsymbol{\Theta}^{\prime}$.
+According to the algorithm, this pair represents the points that are the most promising to be observed next.
+The fifth plot displays the measurements on the quantum computer at the suggested points, indicated by the blue crosses.
 The GP is finally updated using the new measurements—note the change of the blue curve and the shaded area in the plot.
-Finally, the GP’s posterior mean is computed for three equidistant points on the 1D-subspace.
-These points are then used to fit a cosine function according to the functional form of the subspace in Eq.
+Finally, the GP’s posterior mean is computed for three equidistant points on the one-dimensional subspace.
+These points are then used to fit a cosine function, according to the functional form of the subspace given by Eq.
 6.
-This fitted cosine is then used to obtain the analytic minimum of the 1D-subspace.
-We recall that because of the special form of the VQE-kernel and the design of the EMICoRe algorithm, the functional sampled from the GP, i.e, the posterior mean, follows the same functional of Eq.
+This fitted cosine function is then used to obtain the analytic minimum of the one-dimensional subspace.
+We recall that, due to the special form of the VQE-kernel and the design of the EMICoRe algorithm, the function sampled from the GP—i.e, the posterior mean—follows the same functional form as in Eq.
 4.
@@
 4 Numerical Experiments
-In the following, we perform some numerical experiments to demonstrate the performance of EMICoRe in the presence of hardware noise.
-Our experiments focus on the quantum Heisenberg Hamiltonian, a typical Hamiltonian for benchmarking the performance of VQEs [16]
+In the following, we perform numerical experiments to demonstrate the performance of EMICoRe in the presence of simulated hardware noise.
+Our experiments focus on the quantum Heisenberg Hamiltonian, a typical Hamiltonian for benchmarking the performance of VQEs,
 $H=-\left[\sum_{j=1}^{Q-1}\left(J_{X}X_{j}X_{j+1}+J_{Y}Y_{j}Y_{j+1}+J_{Z}Z_{j}Z_{j+1}\right)+\sum_{j=1}^{Q}\left(h_{X}X_{j}+h_{Y}Y_{j}+h_{Z}Z_{j}\right)\right],$ (15)
 where the index of the Pauli matrices $\{X_{j},Y_{j},Z_{j}\}$ refers to the qubit that the Pauli gate is applied to, and open boundary conditions are considered.
-In this work, we chose the couplings to be at the critical point with $(J_{X},J_{Y},J_{Z})=(-1,0,0)$ and $(h_{X},h_{Y},h_{Z})=(0,0,-1)$.
+In this work, we choose the couplings to be at the critical point with $(J_{X},J_{Y},J_{Z})=(-1,0,0)$ and $(h_{X},h_{Y},h_{Z})=(0,0,-1)$.
 The resulting Hamiltonian is also known as the Ising Hamiltonian at criticality.
-In order to assess the performance of the VQE during the optimization, we mainly look at two metrics: energy and fidelity.
+In order to assess the performance of the VQE during optimization, we mainly consider two metrics: energy and fidelity.
 The former is the lowest energy reached by the VQE during the optimization process, i.e., the expectation value of the Hamiltonian with respect to the variational state.
-The fidelity, measures the overlap between the variational (VQE) state and the true ground-state obtained with exact diagonalization.
+The latter measures the overlap between the variational state and the true ground state obtained with exact diagonalization.
+We note that the true ground-state energy $\underline{E}^{\ast}=\langle\psi_{\mathrm{GS}}|H|\psi_{\mathrm{GS}}\rangle$, along with with the corresponding true ground-state wave function $|\psi_{\mathrm{GS}}\rangle$, can be obtained in our case, as the size of the systems we address is sufficiently small.
+This is important as it provides access to a ground truth for benchmarking different algorithms, such as NFT versus EMICoRe.
 We refer to [9] for more details on the metrics we used.
-The ground-state energy $\underline{E}^{\ast}=\langle\psi_{\mathrm{GS}}|H|\psi_{\mathrm{GS}}\rangle$ together with the corresponding ground-state wave function $|\psi_{\mathrm{GS}}\rangle$ can be obtained in our case, because the size of the systems we tackle is sufficiently small.
-This is important as it gives us access to a ground truth for benchmarking different algorithms, i.e., NFT vs EMICoRe.
 To test the stability of the optimization with respect to initialization, we initialize several independent runs with different seeds.
 Specifically, we use $50$ runs for the experiments without hardware noise and $10$ runs for the experiments with hardware noise, see Section 4.1 for more details.
-In all plots we always report the median (50th percentile) as a solid line and the 25th and 75th percentiles being the lower and upper bounds respectively (shaded regions around the solid line).
+In all plots, we report the median (50th percentile) as a solid line, with the 25th and 75th percentiles as the lower and upper bounds, respectively, represented by shaded regions around the solid line.
 The results are plotted as a function of the total number of measurements performed on the quantum computer.
 For all simulations, we use an $L$-layered and fully connected Efficient SU(2) circuit with $D=2(L+1)\cdot Q$ total angular parameters.
-The simulations of the quantum computer are performed with the Qiskit [17] library, while the Python implementation of the EMICoRe algorithm has been adapted from the original code available on GitHub [18].
-All the parameters for EMICoRe are identical to the parameters chosen for the simulation of Fig.
-7 in the original paper [9].
-As noted in Footnote 1, we used $\pi/3$ as the shift for NFT.
-Moreover, in contrast to the original paper [7], we did not include a stabilization step, as our results showed that its use often result in poorer performance.
-We leave further investigation of this phenomenon to future work.
+The classical simulations of the quantum hardware are performed with the Qiskit [16] library, while the Python implementation of the EMICoRe algorithm has been adapted from the original code available on GitHub [17].
+All the parameters for EMICoRe are identical to those chosen for the simulation in Fig.
+7 of the original paper [9].
+As noted in Footnote 1, we used $\alpha=\pi/3$ as the shift for NFT.
+Moreover, in contrast to the original NFT paper [7], we did not include a stabilization step, as our results indicated that its use often led to poorer performance.
+Further investigation of this phenomenon is left to future work.
 (a) Classical simulation of noisy quantum hardware without error mitigation.
@@
 (c) Classical simulation of noisy quantum hardware with ZNE error mitigation.
-Figure 3: Energy (left) and Fidelity (right) for NFT sequential (green) and EMICoRe (red).
+Figure 3: Energy (left) and fidelity (right) for NFT (green) and EMICoRe (red).
 We show results for the critical Ising Hamiltonian using $Q=5$ qubits and $L=3$ layers, for an Efficient SU(2) quantum circuit, and $N_{\textrm{shots}}=1024$ measurement shots.
-The top row shows runs with simulated hardware noise and no error mitigation, while in the second and third rows results for classical simulations of noisy hardware are shown, with TREX and ZNE error mitigation scheme respectively.
-The statistics, e.g., the median (solid line), $25^{\mathrm{th}}$ and $75^{\mathrm{th}}$ percentiles (colored regions surrounding the median), are obtained with $50$ and $10$ independent seeded runs for experiments in Fig.
-3(a) and Fig.
-3(b), respectively.
-A trial density showing the distribution of the runs is shown right next to each plot.
+The top row displays runs with simulated hardware noise and no error mitigation, while the second and third rows present results with TREX and ZNE error mitigation schemes, respectively.
+The median (solid line) and the $25^{\mathrm{th}}$ and $75^{\mathrm{th}}$ percentiles (colored regions surrounding the median) are obtained with $50$ independent seeded runs for experiments in Fig.
+3(a) and with $10$ runs for Figs.
+3(b) and 3(c).
+A density of the trial distribution at the end of the optimization is shown right next to each plot.
 4.1 Hardware Noise Analysis of EMICoRe
-Previous work [9] analyzed EMICoRe’s performance when measurement shot noise is present.
-It was observed, that EMICoRe outperforms NFT more significantly for higher levels of measurement shot noise, i.e., less shots.
-In this proceeding, we investigate the algorithm’s performance when accounting for hardware noise as well.
-In the following, the setup is kept constant for all experiments with $5$-qubits and $3$-layers of the EfficientSU2 circuit and the Ising Hamiltonian with open boundary conditions.
-All simulations are performed for a total of $600$ measurements  [^3] Note that two measurements are performed for each step, i.e., $600$ measurements correspond to $300$ optimization steps.
-and $N=1024$ shots per measurement are used.
+Previous work [9] analyzed EMICoRe’s performance in the presence of measurement shot noise, showing that EMICoRe outperforms NFT more significantly at higher noise levels, i.e., with fewer shots.
+In these proceedings, we extend the analysis to include hardware noise as well.
+In the following experiments, the setup remains constant: we consider a $5$-qubit system with $3$-layers of the EfficientSU2 circuit and the Ising Hamiltonian with open boundary conditions.
+All simulations are conducted with a total of $600$ measurements  [^3] Note that two measurements are performed at each step, meaning that $600$ measurements correspond to $300$ optimization steps., using $N=1024$ shots per measurement.
 In Fig.
-3(a) we run experiments with simulated hardware noise and no error mitigation technique while in Fig.
+3(a), we conduct experiments with simulated hardware noise and no error mitigation.
+In Fig.
 3(b) and Fig.
-3(c) we use twisted readout error extinction (TREX) [19] and zero noise extrapolation (ZNE) [20, 21] as error mitigation strategies.
-The hardware noise is simulated with the Fake5QV1 backend.
+3(c), we apply twisted readout error extinction (TREX) [18] and zero noise extrapolation (ZNE) [19, 20] as error mitigation strategies, respectively.
+Hardware noise is simulated with the Fake5QV1 backend.
 The results in Fig.
-3(a) are obtained by aggregating $50$ independent runs, while Figs.
-3(b) and 3(c) uses only $10$ runs, as the feature was deprecated before more data could be aggregated.
+3(a) are obtained by aggregating $50$ independent runs, whereas Figs.
+3(b) and 3(c) include only $10$ runs, as the feature was deprecated before further data could be collected.
 Table 1 summarizes the results from Fig.
-3, with mean and standard deviations computed over multiple seeded trials, for simulations with and without hardware noise.
-When hardware noise is considered, we observe that EMICoRe significantly outperforms NFT, see Fig.
-3(a), converging faster and more reliably than NFT, e.g., the trial density on the right-hand-side is more peaked around the median.
-We empirically observe that NFT struggles significantly more with hardware noise compared to EMICoRe.
-Notably, neither algorithm achieves the ground-state energy; however, EMICoRe attains a fidelity of nearly $80\%$.
-When using error mitigation, EMICoRe outperforms NFT in both energy minimization and fidelity, being more stable with respect to different initializations, and achieving both lower energies and higher fidelities.
-Table 1: Mean and standard deviation of energy and fidelity calculated at the last optimization step, for all simulations shown in Fig.
+3, reporting the mean and standard deviation computed over multiple seeded trials for simulations with and without error mitigation.
+When hardware noise is taken into account, EMICoRe significantly outperforms NFT, as shown in Fig.
+3(a), demonstrating faster and more reliable convergence.
+For instance, the trial density on the right-hand-side is more peaked around the median.
+These empirical results indicate that NFT is considerably more affected by hardware noise compared to EMICoRe.
+While neither algorithm reaches the ground-state energy, EMICoRe achieves a fidelity of nearly $80\%$.
+With error mitigation, EMICoRe outperforms NFT in both energy minimization and fidelity, demonstrating greater stability across different initializations and achieving lower energies as well as higher fidelities.
+Table 1: Mean and standard deviation of the energy and fidelity at the final optimization step for all simulations presented in Fig.
 3.
-For the energy lower is better, for the fidelity higher is better.
+For the energy, lower values are better; for the fidelity, higher values are better.
 EMICoRe NFT
-Energy $f^{\ast}({\boldsymbol{\hat{\theta}}})$ Fidelity Energy $f^{\ast}({\boldsymbol{\hat{\theta}}})$ Fidelity
+Energy $E^{\ast}({\boldsymbol{\hat{\theta}}})$ Fidelity Energy $E^{\ast}({\boldsymbol{\hat{\theta}}})$ Fidelity
 noise $(-5.472\pm 0.119)$ $(0.782\pm 0.04)$ $(-4.972\pm 0.821)$ $(0.641\pm 0.214)$
@@
 5 Conclusion
-In these proceedings, we leveraged a novel machine-learning-enhanced optimization of VQE algorithms for finding the ground state of the quantum Ising Hamiltonian in the presence of simulated hardware noise.
-This algorithm, termed EMICoRe [9], builds on Gaussian Process Regression, for noise-resilient estimation of the energy, and Bayesian Optimization, to predict the new candidate points to measure at different step in the optimization.
-Our numerical experiments validated the capability of EMICoRe to outperform the state-of-the-art baseline, the NFT algorithm [7], in the presence of simulated hardware noise.
-Even in the presence of simulated hardware noise, EMICoRe is able to converge to lower energies compared to NFT and is more robust with respect to different initializations.
-This robustness highlights the potential of the proposed method for future applications to more complicated quantum systems, and under the challenging conditions of real NISQ devices.
-As a next step, we aim to scale up EMICoRe for larger systems and ultimately evaluate its performance on real quantum hardware.
+In these proceedings, we presented a novel machine-learning-enhanced optimization approach for VQE algorithms to find the ground state of the quantum Ising Hamiltonian in the presence of simulated hardware noise.
+This algorithm, called EMICoRe [9], combines Gaussian Process Regression for noise-resilient energy estimation and Bayesian Optimization to predict new candidate points for measurement at each step of the optimization.
+Our numerical experiments demonstrated that EMICoRe outperforms the state-of-the-art baseline, the NFT algorithm [7], in the presence of simulated hardware noise.
+EMICoRe converges to lower energies than NFT and shows greater robustness across different initializations.
+This robustness underscores the potential of the proposed method for future applications to more complex quantum systems and under the challenging conditions of real NISQ devices.
 Acknowledgments